OpenStudy (yas123):

write an equation in slope intercept form for the line containing the indicated points. (2,4) and (3,5) help !

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

change in y over change in x=slope

OpenStudy (yas123):

can someone solve it

OpenStudy (anonymous):

\[\frac{5-4}{3-2}\]

OpenStudy (anonymous):

not gonna just solve it for you

OpenStudy (yas123):

i 4got the steps on how to solve it thats y i need someone to solve it for me and see the steps on how they do it

OpenStudy (anonymous):

oh... that's different

OpenStudy (anonymous):

u got what I did so far to find the slope?

OpenStudy (yas123):

yea and i got 1 is that right ?

OpenStudy (anonymous):

yep. slope is 1

OpenStudy (anonymous):

so you have y=1x+b pretty much y=x+b

OpenStudy (anonymous):

plug in a point and solve for b

OpenStudy (yas123):

so its not y-4=1(x-2) ?

OpenStudy (anonymous):

4=2+b

OpenStudy (anonymous):

haven't seen it like that... might be correct, but I think that this way is easier.

OpenStudy (anonymous):

so... 2=b y=x+2

OpenStudy (yas123):

ooo oki i see it now thanx

OpenStudy (anonymous):

welcome!